OpenStudy (bluemoon1):

Find the slope of a line parallel to each given line. 1. x-4y = 4 2. 2x + 5y = 0 Medals for fast and accurate answers! Please also show work(:

OpenStudy (steve816):

Remember that a parallel line will have the same slope!

OpenStudy (steve816):

First, you must get the equation into y=mx+b form. Do you think you can do that?

OpenStudy (bluemoon1):

I have no idea how to do this

OpenStudy (bluemoon1):

I don't know how so if you could explain that would be great lol

OpenStudy (bluemoon1):

@steve816

OpenStudy (steve816):

Let's take the first one as an example. This is what you would do: x - 4y = 4 -4y = -x + 4 y = x/4 -1

OpenStudy (steve816):

\[\large y=\frac{ 1 }{ 4}x - 1\]

OpenStudy (steve816):

Therefore, the slope would be 1/4

OpenStudy (bluemoon1):

How about the second one?

OpenStudy (steve816):

Try number 2 on your own!

OpenStudy (bluemoon1):

Okay so I understand number 1 because you are solving for y, right?

OpenStudy (steve816):

You are also solving for y for number 2 too!

OpenStudy (bluemoon1):

2x=-5y x= -5/2?

OpenStudy (bluemoon1):

@steve816

OpenStudy (bluemoon1):

@skullpatrol can you help for last one?

OpenStudy (sshayer):

5y=-2x\[y=-\frac{ 2 }{ 5 }x\] m=?

OpenStudy (bluemoon1):

I don't know how to solve for m

OpenStudy (sshayer):

y=mx+c here c=0 m=?

OpenStudy (bluemoon1):

M=0?

OpenStudy (bluemoon1):

Omg I'm confused

OpenStudy (sshayer):

\[y=-\frac{ 2 }{ 5 }x,y=mx,m=-\frac{ 2 }{ 5 }\]

OpenStudy (bluemoon1):

Oh so that's the answer?

OpenStudy (sshayer):

yes